OpenStudy (anonymous):

For which system of equations is (–2, 2) the solution? A. 3x – 5y = –16 –2x + 3y = 10 B. –4x + 2y = 6 2x – 3y = 10 C. 4x + 3y = 2 5x + 4y = –2 D. 6x + 5y = –2 –8x + 5y = –6

OpenStudy (anonymous):

have u ever heard of the matrix theorem?

OpenStudy (anonymous):

no, never .

OpenStudy (anonymous):

would u like to know because it will make these problems very easy

OpenStudy (anonymous):

yes, thank you.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

how do I do it ?

OpenStudy (anonymous):

The matrix theorem - is an arrangement of numbers of numbers in horizontal rows and vertical columns. This first part will not get the answer to your question but the next part will. Ex. C1 C2 R1 | a c | R2 | b d | Your first step is to cross multiply and then subtract your two answers. a(d) - b(c) Ex. | -1 -2 | | -1 3 | -1(3) - (-2)(-1) -3 - 2 = -5

OpenStudy (anonymous):

ok now the part that will help u solve ur equation is still the matrix theorem but it is a branch of the theorem and is call The Cramers Rule

OpenStudy (anonymous):

it uses the same steps as above just a few more steps added

OpenStudy (anonymous):

Cramers Rule: To solve systems: Ex. 2x + 7y = -3 3x - 8y = -23 you will write the equation into matrix form as so to find your coefficient matrix. | 2 7 | | 3 8 | now u cross multiply and subtract to find ur coefficient. (2) (-8) - (7)(3) -16 - 21 -37 is ur coefficient now u find x pay close attention to how i write this for im running out of time in my class and i cant explain it completely. x= | -3 7 | | -23 -8 | ----------- -37 still cross multilply (-3)(-8) - (7)(-23)